The magnitude and direction of vector, resolve it into its x and y components (2, 45o)
Answer by fractalier(2101)

x = r cos q = 2 cos 45° = 2*(sqrt(2)/2) = sqrt(2)
y = r sin q = 2 sin 45° = 2*(sqrt(2)/2) = sqrt(2)